OpenStudy (anonymous):

Renee took a long flight to visit her family that lives out of the country. She slept for the first half of the trip and watched movies for 1/4 of the remaining flight. If the total trip took 10 1/2 hours, how long did Renee watch movies?

OpenStudy (hoblos):

he total trip took 10 1/2 hours she slept half of it ----> (10 1/2) /2 = 5 1/4 watched movies for 1/4 of the remaining ---> (5 1/4) /4 = 5/4 + 1/16 = (20+1)/16 = 21/16 = 1 5/16 hrs